3. The following information is taken from the balance sheet of Raineer Pumbling:

cash and cash equivalents: $51,000

accounts receivable: $131,000

inventories: ?

total current assets: $290,000

property, plant, and equipment (net): ?

total assets: $510,000

accounts payable: $43,000

note payable (due in two years): $61,000

common stock: $140,000

retained earnings: ?

Determine the missing amounts.
Answer #1

total current assets=cash and cash equivalents+AR+inventory

290,000=51000+131000+inventory

inventory=(290,000-51000-131000)=$108,000

total assets:=total current assets+property, plant, and equipment

property, plant, and equipment=(510000-290,000)=$220,000

Total liabilities=AP+Note payable

=(43000+61000)=$104000

Total assets=Total liabilities+Total equity

510000=104000+(Common stock+Retained earnings)

Retained earnings=(510,000-104000-140,000)

=$266,000